# How structured are the representations in transformer-based vision encoders? An analysis of multi-object representations in vision-language models

### v2

@@ 3 Methods
Dataset: We needed instance segmentation masks to associate the tokens to the object in the image. Hence we used the COCO dataset and created subsets with enough combination of two object categories co-occurring in them. We then excluded images with more than one object category of primary or secondary objects. While choosing the objects across primary and secondary categories we preferred objects more likely to interact in the scene. Within the primary and secondary categories we preferred objects that are similar to each other, so their embedded representations are not naturally distinct, increasing our probe’s sensitivity. We used six task sets with a total of 16,288 images. For example, the first task contains a combination of objects from two sets, i.e. a primary object: animal (cat/dog) and a secondary object: furniture (chair/bench/bed/couch). The dataset and its tasks are detailed in Table 1. We call this set of tasks ’object pair decoding tasks’. The global probes were trained and tested on a larger dataset. We selected the first 40000 images in COCO’s training set for training and the 5000 images in the validation set for testing.
@@
Table 1: For paired object decoding, we use 6 tasks with different number of images in each task. Each task contains images with different variations of objects
Task1 Task2 Task 3 Task4 Task 5 Task 6
@@
Objects 1 cat, dog dining table, person train, bus tv, laptop microwave, oven motorcycle, car
Objects 2 bench, chair, couch, bed pizza, knife, cup, cake traffic light, bench, backpack, handbag mouse, remote, keyboard, cellphone bottle, spoon, knife, cup traffic light, handbag, backpack, bicycle
Linear Probing: We probe the representations in pre-trained networks for our analysis. Probing an information system involves obtaining a representation, usually in the form of a vector from the system in response to an image. Then, we estimate if that particular vector can classify information about the stimuli correctly. The kind of information the probe can learn to classify and the complexity of the probe (i.e. is it just a linear classifier or a complex multilayer NN) indicates the nature of information present in that layer’s representation. We obtain representations from various parts (layers and spatial sections tokens/cells) of the network and use them to understand the network by looking at the classification ability of a linear probe.
Analysed Networks: We analysed the image encoders of seven models: BLIP (ViT-B/16 and ViT-L/14) model for image captioning, CLIP for image-text matching, FLAVA model with an additional multimodal encoder on top of vision and language encoders and finally, VIT-B/16 encoder trained on imagenet21k image for image classification task. We analysed both VIT B/16 and L/14 (Image Transformer) and the Resnet X4 (CNN) image encoder for the CLIP. In CNN analysis, we obtained feature cells instead of tokens, i.e., we used the vector representation of a cell in the feature map by accumulating all the filter outputs at the cell location. In both transformers and CNNs, the place on the feature map representing the object is computed by scaling the object segmentation map to the size of the feature map at the layer. The images are pre-processed with the standard pre-processing function and setting provided along with the pre-trained network instances.
@@ 4 Results
Figure 2: a. Object-pair decoding task results for BLIP: Average decoding performance for different layers (y-axis) and token types (x-axis) over 6 tasks for BLIP. In the subfigures, the y-axis contains variations of where the object-specific tokens (random_obj and avg_obj) are obtained. The different columns show results for 1) decoding the primary object 2) the secondary object 3) the combination of both objects in the image. The decoding pattern remains after averaging, with the tokens from the objects modelling the most useful information for categorising the objects. They are much better than the CLS token, which has to capture the larger scene context. b. Visualisation of cosine similarity to other tokens for a token from primary and secondary object.
@@ 4.1 A snapshot of network’s global and image-specific representation
Decoding the objects in a multi-object scene using the trained model’s representations gives us a picture of how the information about the scene is organised in the network. In our study, we have used two kinds of decoding tasks (object-paired tasks and global decoding tasks) to estimate the architecture-level organisation of these representations. We also generate representation similarity maps to analyse the representation of a single image in a particular model.
Using the paired object decoding tasks, we can see that the objects in the network can be linearly decoded considerably above random accuracy by using a single token representation originating from the object or by using an average of token representations from the object. The general trend followed by these decoding accuracies can be seen in Fig 3. The global object decoding probes, with their 20 class classification setting, provide a good baseline of the ability of the token’s representations to decode between multiple objects. The high correlation of these accuracies for each network (see Fig. 4) to paired-object classification tasks’ results shows that the accuracies are not obtained due to an easy 2-way or 4-way classification task set up in the paired object tasks.
@@
Finally, in Fig. 2, you can see the detailed results for decoding accuracies for the four types of token representation used in our analysis for the paired object decoding tasks. This provides a more detailed global picture of the model’s representations of object pairs from the images. For an instance-wise image-level analysis of representations, we generate similarity maps of the object representations that show how a particular token’s relation to other tokens changes across layers in the model for a particular image. In 2.b, an image shows a representation similarity map for two tokens in an image 1) from the primary object 2) from the secondary object. One can notice how the representations of tokens originating from the same token start having similar representations as we move toward the upper layers. The tokens from the primary object acquire similar representation, but for the secondary object, many tokens that are outside the object also have high cosine similarity in the last layer.
@@
4.2 How different token representations encode objects, their interaction and their importance
@@
In the paired-object tasks, the images consist of a primary object and a secondary object. We see the representation of these two object combinations in each image in four type of token representations from the models. Based on the results from the BLIP model shown in Fig 2 we now discuss the general trend of representations across models. Specific differences are discussed in subsection 4.3.
@@
In our results (see Fig. 4), in all pre-trained models, the primary objects are decoded equally well by the CLS token as the average token representation of the object. There is a decrease in decoding accuracy from primary to secondary object categories. This is partially due to the added complexity of a 4-way classification in the secondary object. However, the CLS token decodes the secondary objects with much less accuracy than both the object-specific tokens (avg_obj and random_obj). The CLS tokens that are optimised for the downstream tasks in each of these networks are expected to model the best information about the scene. But this also means that not all objects are linearly decodable by the CLS token. We observe that the object-specific tokens have better object decoding accuracy as compared to the CLS token (In BLIP Fig 2 primary object CLS:0.94 vs Avg_Obj: 0.97 and secondary object CLS: 0.69 vs Avg_Obj: 0.84). There is a particular degradation in decoding performance using the CLS token for secondary objects, where the accuracy for the CLS token is far below the decoding accuracy of object-specific tokens (avg_obj and random_obj).
@@
Previous analysis showed that the object-specific tokens have the highest accuracy for decoding the particular object from which they originated. Still, they also have high decoding accuracy for the other objects in the image. This accuracy is far above random guess and allows these tokens to decode a combination of objects in the current image (see Fig. 2, for results on BLIP). The final decoding accuracy for the combination is 0.59 when using the primary (avg_obj) object token and 0.69 when using the secondary one. In Fig. 2, one can notice that the primary avg_obj token is much worse at representing the secondary objects than the other way around. These results have implications for the disentanglement of these object-specific tokens. Further, the CLS token also shows a decent decoding accuracy as it is optimised to capture more global information from the image. This accuracy is not higher than the object-specific secondary tokens due to its poor capturing of information about these secondary objects in the scene (CLS token decoding accuracy for the secondary object is 0.69 compared to 0.94 for the primary object, as seen in Fig. 2).
@@
Figure 3: Layer-wise test set decoding accuracy for primary and secondary objects for five pre-trained models in the study. Results for all models are shown in appendix. The accuracies are averaged over the six task sets. In each sub-graph, the y-axis denotes the decoding accuracy and the x-axis denotes the layer at which the accuracy was observed. We observe consistent decoding trends across models with a few variations reported in 4.3.
Figure 4: Results from global image probes: Final layer test set decoding accuracy for 20 object classes for all pre-trained image encoders in the study. We observe that all networks perform way above average using a linear readout for a more difficult 20-class classification task.
4.3 VLM models trained on multi-object images have more modular representations than VIT (trained for single object classification)
@@
In this section, we compare the representations of the models with each other. For the global decoding task, we see that the newer VLMs like FLAVA and BLIP have higher decoding accuracy with their object-specific representations than the other counterparts like CLIP (see Fig 4). Larger models like CLIP-L and BLIP-L counterintuitively perform worse than their base counterparts. Specifically, CLIP-L shows fairly lower decoding accuracy for both single random object and average object token representation (random_obj and Average_obj). In the paired-object decoding tasks, the VLM models follow the overall trend of decoding accuracy being correlated to the results in the global object probing tasks (see3). The relative accuracy trends for various tokens are similar across VLM models. The notable observation is the higher overall decoding performance of FLAVA models, which also shows a higher differentiation of object-wise representations, i.e the random token accuracy is fairly lower in the last layer compared to object-specific tokens.
@@
We observe a difference in the structure of representations due to architecture (Transformer vs CNN) and specific training on multi-object tasks i.e ViT vs other models. Hence, we observe a significant decrease in decoding accuracy using the random CNN unit representation compared to random ViT tokens (In the last layers, primary object ViT: 0.84 vs CNN: 0.72; secondary object ViT: 0.54 vs CNN: 0.45)(see 3). Further, the object-specific tokens in CNNs have lower accuracy while decoding the other objects than their ViT counterpart (In the last layers, primary object ViT: 0.88 vs CNN: 0.8; secondary object ViT: 0.6 vs CNN: 0.59). This indicates that CNN has less entanglement of object-specific information across objects than the ViT counterpart in CLIP. We attribute these results to CNNs not having the ability for the information to travel across units in each layer, hence the object information remains more localised. VIT trained on ImageNet21k shows the least differentiation between object-specific and other tokens as compared to the other Transformer models. Here a random token decoded the object with almost similar accuracy as compared to the CLS token (see Fig. 3). And tokens from one object and similarly decode the category of other objects in the scene. Hence it seems like the scene-level information is more uniformly dispersed in the representations of ViT tokens trained only on single object classification task. This is in comparison to all the ViT models trained on tasks requiring modelling of more than one object for a correct representation of the information required for downstream tasks like text matching, captioning etc. Further, the differentiation of object-specific token and consequently the disentanglement of information is more explicit in these networks than VIT.
@@
4.4 Special tokens may not be special for your downstream task with multi-object images
@@
Figure 5: a. Variation in decoding accuracy between instances of objects ‘in caption’ and ‘not in caption’. Each subplot represents the decoding of the object by its object-specific representation. b. Object detection accuracy on test set using probes trained on three different token representations from CLIP and zero-shot CLIP accuracy.
@@
Following the observations about the lower decoding accuracy of secondary objects using CLS tokens, we wanted to check if the low accuracy results from the object’s importance in the downstream task (captioning in the case of BLIP). Hence, we analysed the decoding accuracy for two sets of data. The set was divided based on whether the object was directly mentioned in the caption generated by the model. Due to this split, we rejected tasks with less than 400 samples in either of the new sets. Therefore, we are left with 3 tasks whose average results are reported in Fig. 5a. In our results in Fig. 5a, we report a decrease in decoding accuracy of the objects not mentioned in captions using all kinds of tokens considered in our analysis (Average final layer accuracy for primary object, which was for ’in caption’: CLS:1 and avg_token: 0.99, drops to CLS: 0.87 and avg_obj 0.93 when the object is ’not in caption’. In the secondary object, the drop is from ’in caption’: CLS: 0.96 and avg_obj: 0.99 to ’not in caption’: CLS: 0.79 and avg_obj: 0.82). This means the network pays more attention to certain objects, and its learning of discriminative features deteriorates for objects not mentioned in the captions. We note that the decrease in the decoding is most pronounced for the CLS token, showing the direct effect of the downstream task on the representation.
@@
We further use this understanding to check for failure in zero shot object detection (i.e. retrieval) tasks using CLIP in a multi-object setting. We use the classification task setting for the global probe. We use the object-specific representation using 1 token and avg_obj token to learn probe to classify the top 20 objects in test set images. Then we evaluate the CLIP zero shot performance to detect objects in the image using the prompt ’An image containing a ’category”. Where the category was replaced with 20 candidate objects. We finally evaluate the accuracy of finding the objects. A fourth multi-class object detection probe is learnt using the CLS token representation. As we compare the accuracy of these four methods on the simple object retrieval task, we find that the probes using the average patch representation gives the best accuracy followed by random_obj token and zero shot model accuracy and multi-class probe (See Fig. 5b).

### v3

@@ 3.4 Dataset
We needed instance segmentation masks to associate the tokens with the object in the image. Hence, we used the COCO dataset and created subsets combining object categories. We then excluded images with more than one object category of primary or secondary type. We also sub-sample to balance the co-occurrence of all object combinations between the primary and secondary categories. This balancing step is important to measure entanglement correctly, as otherwise, the difference in the co-occurrence of primary and secondary objects allows the decoder to learn the presence of one from the other. While choosing the objects across primary and secondary class we preferred objects more likely to interact in the scene. Within the primary and secondary categories, we preferred objects that are similar to each other, so their embedded representations are not naturally distinct, increasing our probe’s sensitivity. We obtain six task sets with a total of 16,288 images. For example, the first task contains a combination of objects from two sets, i.e. a primary object: animal (cat/dog) and a secondary object: furniture (chair/bench/bed/couch). The dataset and its object sets are detailed in Table 1. We call this set of tasks ’object pair decoding tasks’.
For the generalisation task, global probes were trained and tested on a larger dataset of 20 random classes (given in Table 1) beyond the ones used in the main analysis and measures. We selected the first 40000 images in COCO’s training set for training and the 5000 images in the validation set for testing.
@@
Table 1: For paired object decoding, we use 6 object sets with different numbers of images in each set. Each set contains images with different variations of objects. For the global object decoding task, we tested generalisation on 20 randomly chosen objects.
Paired-Object Probe Set 1 Set 2 Set 3 Set 4 Set 5 Set 6
@@
Primary Object cat, dog dining table, person train, bus tv, laptop microwave, oven motorcycle, car
Secondary Object bench, chair, couch, bed pizza, knife, cup, cake traffic light, bench, backpack, handbag mouse, remote, keyboard, cellphone bottle, spoon, knife, cup traffic light, handbag, backpack, bicycle
Global Probe sheep, bear, banana, potted plant, bowl, toilet, horse, apple, fire, parking meter, handbag, snowboard, broccoli, giraffe, stop sign,hydrant, cow, tie, hot dog, truck, wine glass
3.5 Linear Probing
We probe the representations in pre-trained networks for our analysis. Probing an information system involves obtaining a representation, usually in the form of a vector from the system in response to an image. Then, we estimate if that particular vector can classify information about the stimuli correctly. The kind of information the probe can learn to classify, and the complexity of the probe (i.e. is it just a linear classifier or a complex multilayer NN) indicates the nature of information present in that layer’s representation. We do not use a more complex probe because it is difficult to verify if the classification performance for them is a function of better representation or due to the learnt complex relationship in the probe [2, 13]. Hence, we obtain representations from various parts (layers and spatial sections, i.e., tokens) of the network and use them to understand the network by looking at the ability of a linear probe to classify it.
3.6 Analysed models and their configuration
We analysed the image encoders of the nine pre-trained models.
  • Vision-Language Models: BLIP (ViT-B/16 and ViT-L/14) for image captioning, CLIP (ViT-B/16 and ViT-L/14) for image-text matching, and FLAVA with an additional multimodal encoder on top of vision and language encoders.
  • Baseline ViT-B/16: trained on ImageNet21k for image classification.
  • CNN-based CLIP (ResNet X4): an alternative architecture compared to ViT-based CLIP.
  • Self-Supervised Methods: DINO and DINOv2, each trained on images without labels.
In CNN analysis, we obtained feature cells instead of tokens, i.e., we used the vector representation of a cell in the feature map by accumulating all the filter outputs at the cell location. In both transformers and CNNs, the place on the feature map representing the object is computed by scaling the object segmentation map to the size of the feature map at the layer. The images are pre-processed with the standard pre-processing function and setting provided along with the pre-trained network instances.
@@ 4 Results
In this section, we will discuss the results and observations made while looking at the vision encoders through the lens of structured representations. In subsection 1) we see how the images from the paired-object task allow us to see an aggregate and instance-specific view of the representation of various objects in different types of tokens and layers. Subsection 2) shows how the object representations are spread across object-specific tokens and CLS tokens with the example of BLIP. 3) We further discuss the results for all networks, discussing the key variations in the distribution of representations based on the type of network and the training objective. 4) Following up on the result of object-specific tokens having better representations than CLS tokens, we illustrate this effect by relating it to downstream tasks. 5) Finally, we correlate the results obtained till now to unrelated objects in the global decoding task. We see how the measures of structured representation can be used to estimate the layer and network better at retaining object-specific representations in multi-object scenes.
@@ 4.1 A snapshot of network’s global and image-specific representation
Figure 2: a. Paired object decoding task results for BLIP across layers: Average decoding performance for different layers (y-axis) and token types (x-axis) over 6 tasks for BLIP. In the subfigures, the y-axis contains variations of where the object-specific tokens (random_obj and avg_obj) are obtained. The different columns show results for 1) decoding the primary object 2) the secondary object 3) the combination of both objects in the image. The decoding pattern remains after averaging, with the tokens from the objects modelling the most useful information for categorising the objects. The object-specific tokens are much better than the CLS token, which has to capture the larger scene context. b. Visualisation of cosine similarity of highlighted token to other tokens for a token from primary and secondary objects at various layers of BLIP model.
@@
Decoding the objects in a multi-object scene using the trained model’s representations gives us a picture of how the information about the scene is organised in the network. In our study, we have used two kinds of decoding tasks (paired object and global decoding tasks) to estimate the architecture-level organisation of these representations over a set of images. We also generate representation similarity maps to analyse the representation of a single image in a particular model.
@@
Using the paired object decoding tasks, in Fig. 2.a, you can see the detailed results for decoding accuracies for the four types of token representation used in our analysis for the paired object decoding tasks. This provides a more detailed global picture of the model’s representations of object pairs from the images. For an instance-wise image-level analysis of representations, we generate similarity maps of the object representations that show how a particular token’s relation to other tokens changes across layers in the model for a particular image. Fig. 2.b shows the representation similarity map (visualising cosine similarity) for all tokens to two tokens (marked with yellow and blue squares in the original images) in an image 1) to the primary object 2) to the secondary object. One can notice how the representations of tokens originating from the same token start having similar representations as we move toward the upper layers. The tokens from the primary object (dog) acquire similar representation, but for the secondary object (pillow), many tokens outside the object also have high cosine similarity in the last layer.
Finally, we can see that the objects in the network can be linearly decoded considerably above random accuracy by using either a single token representation originating from the object or by using an average of token representations from the object. The general trend followed by these decoding accuracies can be seen in Fig. 3. The global object decoding probes, with their 20 class classification setting with larger train and test sets (40000 and 5000 respectively), provide a solid baseline of the ability of the token’s representations to decode between multiple objects. This setup also tests on completely different sets of objects compared to the paired-object task (Objects in Table 1). The high correlation of these accuracies for each network (see Fig. 5) with paired-object classification tasks’ results (i.e., $M_{1}$ as it uses Avg_Obj accuracy from the paired object decoding task) shows that the accuracies are not obtained due to an easy 2-way or 4-way classification task setup in the paired object tasks, nor by overfitting to a particular set of objects.
@@
4.2 How different token representations encode objects, their interaction, and their importance
@@
In the paired-object tasks, the images consist of a primary object and a secondary object. We see the representation of these two object combinations in each image in four types of token representations from the models. Based on the results from the BLIP model shown in Fig 2, we now discuss the general trend of representations across models. Specific differences are discussed in subsection 4.3.
@@
In our results (see Fig. 4), in all pre-trained models, the primary objects are decoded equally well by the CLS token as the average token representation of the object. There is a decrease in decoding accuracy from primary to secondary object categories, partially due to the added complexity of a 4-way classification for the secondary object. However, the CLS token decodes the secondary objects with notably lower accuracy than both the object-specific tokens (avg_obj and random_obj). The CLS tokens, optimised for the downstream tasks in each network, are expected to model the best information about the scene. Yet this also means that not all objects are linearly decodable by the CLS token. We observe that the object-specific tokens have better object decoding accuracy compared to the CLS token (in BLIP, Fig. 2, the primary object has CLS: 0.91 vs Avg_Obj: 0.92 and the secondary object has CLS: 0.66 vs Avg_Obj: 0.81). There is a particular degradation in decoding performance using the CLS token for secondary objects, where its accuracy is far below the decoding accuracy of object-specific tokens (avg_obj and random_obj).
@@
Previous analyses showed that object-specific tokens have the highest accuracy for decoding the particular object from which they originated. Still, they also have high decoding accuracy for the other objects in the image. This accuracy is far above random guess and allows these tokens to decode a combination of objects in the current image (see Fig. 2, for BLIP results). For example, the final decoding accuracy for the combination is 0.57, when using the primary (avg_obj) object token and 0.65 when using the secondary one. In Fig. 2, one can notice that the primary avg_obj token is much worse at representing the secondary objects than the other way around. These findings have implications for the disentanglement in object-specific tokens. Further, the CLS token still shows a decent decoding accuracy as it is optimised to capture more global information from the image. However, it remains lower than the object-specific secondary tokens, due to its weaker capturing of information about these secondary objects in the scene (CLS token decoding accuracy for the secondary object is 0.66 compared to 0.91 for the primary object, as seen in Fig.2).
@@
Figure 3: Layer-wise test set decoding accuracy for primary and secondary objects for pre-trained models in the study. Results for all models are shown in the appendix. The accuracies are averaged over the six object sets. In each sub-graph, the y-axis denotes the decoding accuracy, and the x-axis denotes the layer at which the accuracy was observed. We observe consistent decoding trends across models with a few variations reported in Section 4.3.
4.3 Variation of representations across models with different training objectives and architectures: key insights
@@
In this section, we compare the representations of the models with each other. Overall, across the models, the decoding accuracy of the object-specific tokens of the paired-object decoding tasks correlates with their global decoding task performance (see Fig. 5). The relative accuracy trends for various tokens are similar across most VLM models. A notable observation is the higher overall decoding performance of FLAVA, BLIP and BLIP-Large models, which also shows a higher segregation of object-wise representations (i.e., the random token accuracy is fairly lower in the last layer compared to object-specific tokens). These models also score high on our object binding measure and relatively low on entanglement (except BLIP-L and CLIP-L). The two self-supervised models, DINO and DINOv2, show a slightly lower score on object binding for secondary objects, yet their representations are more disentangled, which allows them to model the background objects relatively better. Interestingly, we observe in Fig.3 that in DINOv2, the decoding accuracy of random_obj is higher than other models, aligning with its reported performance on use for semantic segmentation tasks using token-level classification.
@@
We observe a difference in representations’ structure due to architecture (Transformer vs CNN) and specific training on multi-object tasks (i.e., ViT vs other models). Consequently, there is a significant decrease in decoding accuracy using the random CNN unit representation compared to random ViT tokens (in the last layers, primary object: ViT 0.84 vs CNN 0.72; secondary object: ViT 0.54 vs CNN 0.45; see Fig. 3). Further, the object-specific tokens in CNNs have lower accuracy while decoding the other objects in the scene than their ViT counterpart (in the last layers, primary object: ViT 0.88 vs CNN 0.8; secondary object: ViT 0.6 vs CNN 0.59). This indicates that CNN has less entanglement of object-specific information across objects than the ViT counterpart in CLIP. We attribute these results to CNNs not having the ability for the information to travel across units in each layer; hence, the object information remains more localised. ViT trained on ImageNet21k shows the least differentiation between object-specific (Random_Obj: 0.88) and other tokens (Random: 0.84) compared to the other Transformer models. Here, a random token decodes the object with almost similar accuracy as the CLS token (CLS: 0.88; see Fig. 3), and tokens from one object can similarly decode the category of other objects in the scene. Hence, the scene-level information appears more uniformly dispersed in the representations of ViT tokens trained only on single object classification tasks. We note that this difference can be because the other models have been trained on objectives that require correct modelling and representation of multiple objects for downstream tasks like text matching, captioning, etc. Hence, the differentiation of object-specific tokens from random tokens and, consequently, the segregation of information is more explicit in the last layer in many of these networks than in ViT(trained for single object classification).
Turning to the global decoding task, we see that the newer VLMs like FLAVA and BLIP have higher accuracy for their object-specific representations than counterparts like CLIP (see Fig 5). A larger model (BLIP-L) performs better than its base version. However, the layer-wise analysis (see Fig. 3) indicates that the CLS token outperforms object-specific tokens for CLIP-L. This points to a different binding mechanism, where the object-specific tokens have low classification accuracy across object categories, but the CLS token effectively binds both primary and secondary objects in the higher layers. Consequently, one can note that CLIP-L has the best scene-level decoding accuracy (0.66) when using the CLS token to decode the combination of primary and secondary objects. On the other hand, ViT trained on single object classification has the poorest accuracy in decoding a combination of objects using the CLS token (0.49). However, its object-specific regions maintain fairly good representations for individual objects (Avg_Obj $A_{pp}$= 0.93 and $A_{ss}$=0.80).
Lastly, the decoding accuracy for object-specific tokens degrades the most on secondary objects, which is also correlated with its lower accuracy on objects ’not mentioned in the caption’ in Fig. 5. Our layer-wise analysis in Fig.3 also shows that in BLIP-L and many networks like ViT, DINOv2, CLIP, and BLIP, the secondary object accuracy starts decreasing in the final few layers, showing the effect of the training objective. We explore that effect further in the next two sections.
@@
4.4 Special token may not be special for your downstream task with multi-object images
@@
Figure 4: a. Variation in decoding accuracy between instances of objects ‘in caption’ and ‘not in caption’. Each subplot represents the decoding of the object by its object-specific representation. b. Object detection accuracy on the secondary objects in paired using probes trained on three different token representations from CLIP and zero-shot CLIP accuracy.
@@
Following up on the observations about the lower decoding accuracy of secondary objects for CLS tokens, we wanted to check if the low accuracy results are related to the object’s importance in the downstream task (captioning in the case of BLIP). Hence, we analysed the decoding accuracy for two sets of data. We divided the data based on whether the object was directly mentioned in the caption generated by the model. Due to this split, we rejected object sets with less than 400 samples in either of the new sets. Therefore, we are left with 3 object sets whose average results are reported in Fig. 4.a. From Fig. 4.a, one can see a decrease in decoding accuracy of the objects not mentioned in captions using all kinds of tokens considered in our analysis (average final layer accuracy for primary object, which was ’Object in caption’: CLS: 1 and avg_token: 0.99, dropping to CLS: 0.87 and avg_obj 0.93 when the object is ’not in caption’. In the secondary object, the drop is from ’Object in caption’: CLS: 0.96 and avg_obj: 0.99 to ’not in caption’: CLS: 0.79 and avg_obj: 0.82). This means the network pays more attention to certain objects, and its learning of discriminative features deteriorates for objects not mentioned in the captions. We note that the decrease in decoding accuracy is most pronounced for the CLS token, showing the direct effect of the downstream task on the representation.
@@
Using CLIP in a multi-object setting, we further use this understanding to check for failure in zero-shot object detection tasks. We use the classification task setting of the secondary objects in the paired-object probe. Specifically, the object-specific representation (random_token and avg_obj token) is used to learn probes to classify the secondary objects in images. We also evaluate the CLIP zero-shot performance to detect objects in the image with the prompt "An image containing a ’category’", where the category was replaced with 4 object categories of the set. We finally evaluate the accuracy of detecting the objects. The results confirm the original CLIP paper by showing that learning a probe gives better accuracy than zero-shot accuracy in CLIP (Probe CLS: 0.65 vs Zero Shot: 57; see Fig. 4.b). Further, we also note that the object-specific regions in the token space achieve far better representations than the CLS token (Random_Obj: 0.76 and Avg_Obj: 0.77). This emphasises that even though the CLS token is used to classify the objects in practice, comparing it to representations from the token space, we find that the object-specific regions perform much better in classification for CLIP.
4.5 Structured representations in the network are better to retain the representation of objects in the background
Figure 5: The figure shows both measures of modularity evaluated for all vision encoders. Along with this, the figure has the accuracy of global probing task for avg_obj token evaluated for ’In caption’: Objects mentioned in the caption and ’Not in caption’: Objects not mentioned in the caption. The results for each network are from the layer with the best $M_{1}$ across the layers. We see how the first measure of modularity $M_{1}$ correlates highly with the accuracy of decoding background tokens across networks (Pearson’s correlation coefficient = 0.94, p = 0.001). This also shows that the measures from the paired object task are highly predictive of general representation of objects in the network. The second measure of modularity (Measure of entanglement) $M_{2}$ correlates (Pearson’s coefficient = 0.74, p = 0.02) with the drop in accuracy for objects ’not in the caption’ compared to the objects mentioned ’in the caption’. Hence, more modular representations favour better representation of individual background objects.
Table 2: Table showing the predicted layer with maximum object binding measure ($M_{1}$) and the actual layer with the best representation of background objects, i.e. Global Probe:’not in the caption’ in the global decoding task.
Model Layer Predicted Best layer
DINO 8 7
DINO2 10 10
CLIP_Large 23 23
BLIP_Large 20 21
CLIP 10 9
BLIP 9 9
FLAVA 9 8
VIT 7 10
CLIP ResnetX4 3 3
As discussed in the previous section, the object-specific tokens in the main token space of the network have better ability to retain and model object-specific information for most objects than the CLS token. This makes them especially useful for tasks requiring information about objects that may not be the scene’s focus (according to the training distribution). However, a key question remains: how do we quantify this ability of object-specific tokens across layers and networks? Our proposed measures of structured representations address this by evaluating the degree to which networks model and maintain separate object-wise information in the token space. In Fig 5, we show these measures across all networks using the Avg_Obj tokens. Specifically, $M_{1}$ measures the decoding accuracy of the secondary object $O_{s}$ using secondary object token $T_{s}$. We find (see Fig 5) that when networks exhibit strong modular representations of secondary objects, they also decode other 20 "background" objects more effectively if those objects are not mentioned in the COCO captions. Indeed, there is a strong Pearson’s correlation of +0.94 with the object binding measure $M_{1}$ and a moderate inverse correlation of -0.53 (p-value = 0.13) with the entanglement measure $M_{2}$. The role of structured representations becomes clearer when comparing ’not in caption’ objects to ’in caption’ objects: the token entanglement measure $M_{2}$ correlates 0.74 (p-value = 0.02) with the drop in accuracy for ’not in caption’ objects. This indicates that the more segregated the information in the tokens, the better a network’s object-specific representation can be in equally retaining and decoding objects not mentioned in the caption (background objects). Hence, this shows that more modular representations in vision encoders, i.e. representation agreeing to our two properties of 1) binding and 2) segregation, are signs of better retention of object-specific representation of even less important ’objects not in the caption’.
Our analysis also shows that in most encoders, the last layer is not the best choice for decoding the objects using the object-specific tokens from the networks; in models such as BLIP-L, object-wise representation degrades in the last two layers (0.87 in layer 20 to 0.73 in layer 23 for secondary objects), possibly due to the pressure of the objective function. In Table 2, we compare the estimated choice of the layer using the object binding measure $M_{1}$ and the layer with the best accuracy on background objects (’not mentioned in the caption’). Further, a lower entanglement score helped guide the choice when object binding measures are similar across layers. We see a good relative match between the estimated best layers using measures of modularity obtained using paired-object tasks and the best layers for the larger 20-class global decoding task. Hence, our measures of structured representations can be helpful in ascertaining which network and particular layers are useful to train decoders for tasks requiring information about background objects in multi-object scenarios.